Eng Half Night Plan UT 26 May 2019

  1. TCS2019A build - no specific tests needed
    • switch it on during day (after handover), so we can troubleshoot possible problems
    • leave it during night unless new SW problems occurs
  2. DONE AGW-1 On-sky checkout - 0.5-1hr
  3. DONE LUCI-1 Rotator Center Check - 0.5 hr
    • N30 fieldstop, N375 Camera, track mode, 4 PAs (0,180,90,270)
    • Collimate first with an active preset then resend preset in track mode
    • Melottle 111 (Coma Bernices) RA 12 Dec +26
    • Files located at: /home/observer/Service_Observing/E_Tasks_May2019/LUCI_Tasks/
  4. DONE LUCI-1 zeropoint rotator measurement - 0.5 hr - LUCI-1 only (no masks in LUCI-2) - Attempted - not confident in 2nd mask
    • Above 30 elevation all night - should do both masks to get enough alignment boxes
      • ID 942616 luci1.8.magmask1.lms 17:56:59 +65 18 42
      • ID 924246 luci1.874.CaT3.lms 14:19:35 +52:49:29 - high elevation first half
      • Files located at: /home/observer/Service_Observing/E_Tasks_May2019/LUCI_Tasks/E_LUCI_1_ZP_Test
  5. LUCI-1/AGW-1 transform data collection (take bino data - DX for "free") - 2.5 hours (update PCSconfig with new rotator ZP then take transform - replace old ZP afterwards)
  6. Collimation Check on-sky for MODS (If MODS-2 IMCS lasers replaced) - Postponed
  7. Shivaei Program - 2hrs (If MODS-2 IMCS lasers replaced) - Postponed
  8. Poor weather backup program - any amount of time - targets cover entire night
    • H-band library - K and M stars
    • H-band library - case example of QSO rest-frame 1.6micron CO-bandhead - filler
    • Located at: /home/observer/Service_Observing/E_Poor_Weather_HbandLibrary (remote room)


Observing Log UT 26 May 2019

OSA: J. Williams

ISp: G. Bechetti

ISA: B. Rothberg

Software: P. Kubanek

AO: D. Miller

UT 01:49 - minor hiccup, TCS was not mounted on OBS1/telescope - so I wasn't sure about where to access the config files. I can edit the files directly on the TCS computer - this is the new policy (now I know!)

UT 01:50 - moving telescope to zenith and park for LUCIs, will init all

-NOTE: configuration files now need to be accessed on TCS computer directly

BGTech needs to be updated - looking for conf files...

UT 02:22 - opening enclosure

For hotspot (AGW1/2/3/4)

on oac (must login as oac@oac):

/lbt/oacontrol/current/etc/oacontrol.conf

images should be located (on obs machine - logged in as lbto)

/lbt/data/share/tcs/lbto.201N.XX.YY/

UT 02:38 - taking two 2.51sec dark test exposrures

luci1.201905026.0009-0010

luci2.201905026.0026-0027

UT 02:41 - sending pointing preset

UT 02:48 - starting hotspot measurement

-guiding stopped

ptincrement 0 CA -3

-setuxy -u 1 -x 0 -y 612.5

- first exposure will fail so do
gcsclient readGuideCam left -e 1000 (1 sec)

then

gcsclient readGuideCam left -e 90000

-prompt isn't returning

UT 02:56 - running 90sec again, then after 2-3sec hit slew to hold 360

left_guider000002.fits

WAY too bright....blasted the detector

-moving to a 12th or 13th mag star for collimation

-need to take a longer exposure because rotator is now slower than before

-There appears to be a fiber on in the middle of the secondary

UT 03:06 -Trying this again:

left_guider000003.fits

ptincrement 0 CA 3

using 180seconds

-measuring center X=117 Y=112 multiply by 2 ==> X=234 Y=224

DETSEC [16:536,40:508] X+15 , Y + 39

X= 249 , Y= 263

gcsclient getHotspot left returns X= 251 Y=256

1 pixel = 1/20 arcsec for Potsdam AGw; 0.600 mm per arcsec

move +2 pixels in X, +7 pixels in Y - J.Hill suggests given the wobbly look of the circle, we are ignoring the X offset

+0.21 mm

From wiki:

Decreasing RAD ooffset moves the star image (center of rotation) toward larger Y pixel values.

So I want to move towards smaller Y so we will go +210 microns

current value -417912, new value is: -417702

in OAC terminal: rdwrconfig -u 1

UT 03:23 - restart GCS left

UT 03:24 - resending preset to same star

-now to verify

guiding stopped, sent to 360 rotation

ptincrement 0 CA -3

setxy -u 1 -x 0 -y 612.5 - seems to be hanging

type "getxy -u 1" to see where we are -- we are close

UT 03:36 - sending command again...hanging

-CTRL C'd - proceeding with image

-seems to be taking a lot longer than 180seconds...

UT 03:45 -log shows an error resolving OBS2 network issue

UT 03:47 - AZCAM server timed out

-tested 1 seconds image from OBS4 left_guider000004.fits - it worked

-trying again but taking image via OBS4

left_guider000005.fits

-measuring center X=108 Y= 117 multiply by 2 ==> X=216 Y=234

DETSEC [16:536,40:508] X+15 , Y + 39

X=231 , Y= 263

gcsclient getHotspot left returns X= 251 Y=256

1 pixel = 1/20 arcsec for Potsdam AGw; 0.600 mm per arcsec

now 20pixels off in X and Y didn't change from before

-Consensus is to revert to original values - first meaured offset was a few pixels, the update caused a 20pixel shift in X

UT 03:57 - restarting GCS

UT 03:58 - reauthorizing for binocular LUCI

UT 04:03 - moving to Rotator Center tests - Binocular LUCI

UT 04:14 - sending active preset to Melotte 111 (Coma Bernices)

MOS Error - LUCI-2

El 78, Rotator 311

failed at FPU to turnout

-grabbing the mask - PICKER ARMS

-error while returning mask 3 to cabinet

While tracking this down colimation on left exploded

-8000 steps 1.1 mos mcu - moving to neg limit - failed to grab

SG off picker arms open

robot is in right position

check preconditions hold to locked, then execute

picker arms open, mask locked in fpu

initialize mos

turn off hc on rotator, init tirned off hc on spindle

initializing again

UT 04:39 - MOS error recovered

L1 0011 test exposure active preset

L2 0028.0029 test exposure active preset - field stop is off - moving

UT 04:48 - starting rotator center tests

L1 0012 off, 0013 on PA 0

L2 0030 off, 0031 on PA 0

--- don't do off exposures in the sequence just take an "on" in each PA and use the "off" from the active preset (first exposure in sequence for active preset)

-IGNORE LUCI images L1 0011-0013, L2 0028-0031

-scripting is not working

UT 05:02 - sending a track preset to target with PA=0

- probe is actiing up, couldn't move probe on Track preset, homing probe now

- taking a long time to home

UT 05:10 - probe is homed

-IT 7631 - relaxing pmac position tolerance

pmacsend -u1 128 - returns nothing on oac except a timestamp - its 'i' not 'l'

UT 05:13 - J.Williams will home AGW again

- sending track preset - failed AGW-1 did not home

-restart oacontrol, restart gcs

-logged back into oac - send command "pmacsend -u 1 i128' - returned an error connecting to AGW-1, restarting GCS, waiting...

-after gcsrestart the gui shows an error - failed to communicate with umac

-still can't connect umacs after GCS, calling P. Kubanek

-connected and working

-changing pmacsend to 48

UT 05:35 - pmacsend -u1 'I128=48'
UT 05:35 - sending track preset again

- will do an acquire manually to center up the star then proceed with Track scripts

acquire preset allows a bigger pointing change

gcsclient setGSsfp left - this center probe

gcsclient setGSsfp right - ""

gcsclient startAcquire left - takes a look of images to find start - OSA can then center star on both side

gcsclient startAcquire right - ""

- can take the PA = 0 image

luci1.20190526.0016 - ignore - guideprobe is on axis

luci2.20190526.0034 - ignore guideprobe is on axis

Rotation Center Data

luci1.20190526.0017 - PA 0

luci2.20190526.0035 - PA 0

luci1.20190526.0018 - PA 180

luci2.20190526.0036 - PA 180

luci1.20190526.0019 - PA 90

luci2.20190526.0037 - PA 90

luci1.20190526.0020- PA 270

luci2.20190526.0038 - PA 270

Taking sky offs: - moved +60" in DETX sync offset - according to OT this is blank

luci1.20190526.0021

luci2.20190526.0039

UT 05:58 - authorizing LUCI-1 only (blinds in LUCI-2)

Rotator ZP test - use two MOS masks to acquire and measure rotation needed

CaT3 ID 924246

UT 06:00 - sendind acquisition script

Using 4/5 stars (30, 31, 33, 34)

X-shift 1.269"

Y-shift -0.281"

Angle: 0.372 (initial guess angle 0.291)

Err RA 0.09 Dec 0.072

Moving to next mask ID 942515 magmask1

UT 06:24 - acquiring

UT 06:30 -collimation is really struggling

-initial angle guess 0.188

- mask is showing 30-39 alignment stars - but only 4 are showing up (30, 31, 35, and 37)

Xshift -0.08

Yshift -0.148

Angle 0.670

Err RA DEC 0.191 0.348

Angle seems WAY too large

UT 06:44 - AGW-1 error failed to move, offset was only -0.09" -0.15" - error

I don't see stars in boxes - could be the WFS running away and blurring out stars - or the mask/alignment stars are bad - Need to do a rotator zp using another method

UT 06:48 - P.Kubanek trying to restart OAC

UT 06:49 - handing over LN - too many errors, issues, and not enough time to start (or even confident we should) agw transform

-- BarryRothberg - 25 May 2019
Topic revision: r4 - 26 May 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback